OpenStudy (peterg2001):

help will medal Solve the following equation, and check your solution. Be sure to show all work in order to receive full credit. Solve 24 - 3x = -27.

OpenStudy (anonymous):

x=17

OpenStudy (anonymous):

24 - 3x = -27 -24 - 24 -3x = -51 /-3 /-3 x = 17